# Text Generation Beyond Discrete Token Sampling

### v1

@@ 5.3 Baselines
To quantify the benefit of MoI, we compare it with two decoding schemes that keep the underlying model architecture and sampling mechanism fixed. The primary baseline (Standard) is the widely used nucleus sampling with temperature scaling [1]. It represents the default inference recipe shipped with each model. Our second baseline (Direct Mixture) constructs the input representation as a simple weighted sum of token embeddings using the softmax probabilities as coefficients, i.e., computing the value of $\boldsymbol{h}_{t}$ as $\sum_{i=1}^{V}p_{t,i}\boldsymbol{e}_{i}$. Unlike MoI, it performs no Bayesian reconciliation between the distribution and the sampled token, providing a stringent ablation for assessing the value of our posterior estimator.
@@ 6.2 Hyperparameter Importance Analysis
Fig. 2 provides two complementary perspectives on hyperparameter importance. The left plot tracks expected performance gain when optimizing each parameter individually through best-of-N-shots tuning, with experiment setup explained in Section B.1. As N increases from 1 to 15, $\beta$ consistently yields the highest gains, reaching nearly 7.5% improvement at N=15, while top-p and temperature plateau at approximately 6.0-6.5%. This separation becomes particularly pronounced after N=10, suggesting that $\beta$’s impact grows with more extensive search.
The right panel quantifies each parameter’s importance through random forest regression analysis, with experiment setup explained in Section B.2. With inputs as hyperparameters and accuracy as the target, this reveals $\beta$ as the dominant factor (importance score of 0.41), followed by top-p (0.32) and temperature (0.27).
@@ Appendix A Comparing over Hyperparameter Grid Search
Appendix B Additional Setups
B.1 Best-of-N Analysis Setup
@@
B.2 Random-Forest Regression Analysis Setup
@@
B.3 Setup for Case Study
@@ Appendix C Hyperparameters
Appendix D Implementation Details

### v2

@@ 5.3 Baselines
To quantify the benefit of MoI, we compare it with two decoding schemes that keep the underlying model architecture and sampling mechanism fixed. The primary baseline (Standard) is the widely used nucleus sampling with temperature scaling [1]. It represents the default inference recipe shipped with each model. Our second baseline (Direct Mixture) constructs the input representation as a simple weighted sum of token embeddings using the softmax probabilities as coefficients, i.e., computing the value of $\boldsymbol{h}_{t}$ as $\sum_{i=1}^{V}p_{t,i}\boldsymbol{e}_{i}$. Unlike MoI, it performs no Bayesian reconciliation between the distribution and the sampled token, providing a stringent ablation for assessing the value of our posterior estimator. We also tried directly feeding the mixed output hidden states, but we found that the models cannot make sense of the hidden states without retraining.
@@ 6.2 Hyperparameter Importance Analysis
Fig. 2 provides two complementary perspectives on hyperparameter importance. The left plot tracks expected performance gain when optimizing each parameter individually through best-of-N-shots tuning, with experiment setup explained in Section C.1. As N increases from 1 to 15, $\beta$ consistently yields the highest gains, reaching nearly 7.5% improvement at N=15, while top-p and temperature plateau at approximately 6.0-6.5%. This separation becomes particularly pronounced after N=10, suggesting that $\beta$’s impact grows with more extensive search.
The right panel quantifies each parameter’s importance through random forest regression analysis, with experiment setup explained in Section C.2. With inputs as hyperparameters and accuracy as the target, this reveals $\beta$ as the dominant factor (importance score of 0.41), followed by top-p (0.32) and temperature (0.27).
@@ Appendix A Comparing over Hyperparameter Grid Search
Appendix B Generalization of a single hyper-parameter setting across tasks
Table 4: Additional results on four benchmarks with four large language models. For every (method, model) we tune the decoding hyperparameters on AIME and CountDown 4, fix the best setting, and reuse it when assessing GPQA-Diamond and LiveCodeBench. Accuracy (%) is reported on AIME, Count Down 4, GPQA-Diamond, and pass@1 is used on LiveCodeBench. Standard uses conventional sampling, that is temperature–scaled nucleus sampling and MoI is our full approach.
Model Method Input Info. AIME CountDown4 GPQA-D LiveCodeBench Avg
QwQ-32B Standard Output Token 76.89 78.04 58.08 76.48 72.37
MoI 78.44 79.22 61.41 77.85 74.23
   Gain vs. Standard Token + Dist. +1.55 +1.18 +3.33 +1.37 +1.86
Nemotron-Super-49B Standard Output Token 54.67 56.77 63.64 40.35 53.86
MoI 57.11 59.22 64.95 40.90 55.55
   Gain vs. Standard Token + Dist. +2.44 +2.45 +1.31 +0.55 +1.69
Gemma-3-27B Standard Output Token 24.44 55.21 45.45 32.37 39.37
MoI 26.67 57.19 46.26 32.92 40.76
   Gain vs. Standard Token + Dist. +2.23 +1.98 +0.81 +0.55 +1.39
DAPO-Qwen-32B Standard Output Token 64.67 70.31 42.93 56.91 58.70
MoI 61.56 75.00 43.94 57.50 59.50
   Gain vs. Standard Token + Dist. -3.11 +4.69 +1.01 +0.59 +0.80
A practical concern when adding new decoding hyperparameters is whether the values tuned on one task will transfer to others. To investigate this, we conduct a grid search over the hyperparameters described in Section 5.4. For each (method, model) pair, we keep the configuration that maximizes the average accuracy on AIME & CountDown 4 and freeze it for the remainder of the study.
Table 4 shows the results on the held-out GPQA-Diamond and LiveCodeBench. The single-tuned setting consistently outperforms Standard nucleus sampling on 15 of the 16 (model, task) pairs.
Overall, these results suggest that MoI requires only modest tuning effort: once the hyperparameters are calibrated on a small proxy set, they generalize robustly to new tasks and domains without much further adjustment.
Appendix C Additional Setups
C.1 Best-of-N Analysis Setup
@@
C.2 Random-Forest Regression Analysis Setup
@@
C.3 Setup for Case Study
@@ Appendix D Hyperparameters
Appendix E Implementation Details